OpenStudy (anonymous):

How would I use the definition of the derivative to find (2/(sqrt(x))' ?

OpenStudy (anonymous):

definition of a derivative, from first principles: \[df/dx = \lim_{h \rightarrow 0}[f(x+h) - f(x)]/h\]\[d(2x^{-1/2})/dx = \lim_{h \rightarrow 0}(1/h)(1/\sqrt{x+h} - 1/\sqrt x)\] next step is to add the fraction when you add the fraction, rationalise the NUMERATOR by multiplying by the conugate the x's drop out, then the h cancels with 1/h & it's downhill from there If you need more explicit help, say so. The trick is rationalising the numerator to be able to cancel out h's

OpenStudy (anonymous):

typos = multiply by the CONJUGATE

OpenStudy (anonymous):

Jashcache do you want me to write it out? or do you follow his comment?

OpenStudy (anonymous):